OpenStudy (boom):

9^2n less than equal to 1/6561

OpenStudy (lgbasallote):

9 = 3^2 6561 = 3^8 \[\large 3^{4n} \le 3^{-8}\]

OpenStudy (lgbasallote):

\[\large 4n \le -8\] \[\large n \le -2\]